A loan's quirky terms require the first payment to be made at the moment the money is lent. Payments are $65 every month, the monthly interest rate is 0.44%, and the loan is paid off in equal payments over 6 years. How much money was borrowed? Answer and round to the nearest cent.
Answer #1

Payment is in form of Annuity due

PMT = -65
Number of Periods (n) = 12*6 = 72
Interest per month = 0.44%

Amount of Money Borrowed =(1+r) * PMT *(1-(1+r)-n)/r = (1+0.44%) * 65 * (1-(1+0.44%)-72)/0.44% = 4021.28

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well